HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE ROTATIONAL SPECTRUM OF THE METHANOL ISOTOPOLOG CH_{3}^{17}OH UP TO 1.1 THz

\underline{H. S. P. MÜLLER} , F. LEWEN, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; V. V. ILYUSHIN, Institute of Radio

Astronomy of NASU, Mystetstv 4, 61002 Kharkiv, Ukraine

Numerous isotopic species of methanol have been detected in the warm parts of star-forming regions. We have started a program to investigate their rotational spectra to facilitate searches for these in space and, in the longer run, to tackle the issue of torsion-vibration interaction that has hampered quantitative analyses of CH_{3}OH and its isotopologs beyond v_{t} = 2.

After CD_{3}OH,^{1} CD_{3}OD (submitted), and CH_{3}OD (in preparation), we embarked on a study of the rotational spectrum of CH_{3}^{17}OH employing a sample enriched in 17O to 20%. The^{17}O quadrupole splitting is usually quite well resolved at low values of J. Asymmetric broadening through this splitting can affect transitions with high rotational quantum numbers and at high frequencies. The assignments were comparatively straightforward despite the very limited previous data in the literature.^{2} Thus far, we have assigned extensively a- and b-type transitions in v_{t} ≤ 1 and got a fit with weighted rms deviation of 1.1 for 1540 measured line frequencies with J ≤ 40. As in the previous cases, we expect perturbations in v_{t} = 2 through torsion-torsion interaction with higher torsional states to prevent a global treatment of v_{t} ≤ 2 within experimental uncertainties. The analysis is carried out applying the rho axis method and the RAM36hf program code.^{3}

The work in Cologne was supported by the Deutsche Forschungsgemeinschaft (DFG) via the collaborative research center SFB 956, project area B3, and through the Gerätezentrum “Cologne Center for Terahertz Spectroscopy”. The work in Kharkiv was supported by the Volkswagen Stiftung. The assistance of the Science and Technology Center in the Ukraine is acknowledged (STCU partner project P756).

1doi: 10.1051/0004-6361/202142326, V. Ilyushin, H. S. P. Müller, J. K. Jørgensen, et al., Astron.Astrophys. 658 (2022) A127

2doi: 10.1016/0022-2852(91)90405-Y, Y. Hoshino, M. Ohishi, K. Takagi, J. Mol. Spectrosc. 148 (1991) 506-508

3doi: 10.1016/j.molstruc.2020.128037, R. Motiyenko, V. Ilyushin, J. Demaison, et al., J. Mol.

Struct. 1213 (2020) 128037 p-number: p108  Submitted on Wed, 14 Jun 2023 09:49:18 +0200